OpenStudy (anonymous):

cos^-1(cos4pi/5) can someone help me with this one? And explain how they got the answer as well Will fan and medal

OpenStudy (anonymous):

Is cos^-1 arccosine (the inverse of the function) or 1/cos?

OpenStudy (anonymous):

the inverse

OpenStudy (anonymous):

That makes things nice ^_^ So the inverse of cosine does the opposite operation that cosine performs on a number. So whereas \[cos \left( \frac{4 \pi}{5} \right) \approx -.809\] if you perform an inverse function on a function, you just get the original input, so \[ cos^{-1}\left( cos\left( \frac{4 \pi}{5} \right) \right) = \frac{4 \pi}{5}\] They kinda cancel each other out.

OpenStudy (anonymous):

If there's a negative in the parentheses does that change things? @AllTehMaffs

OpenStudy (anonymous):

? like a negative before the cosine? \[cos^{-1}\left( -cos \left( \frac{4 \pi}{5} \right) \right) \] ?

OpenStudy (anonymous):

or before the -4pi/5 ?

OpenStudy (anonymous):

Like for number 42. on my homework cos^-1 [cos(-5pi/3)}

OpenStudy (anonymous):

The output of the cos^-1 is always what's in the argument of the cosine that it's inverting, so the answer to that one would be (-5 pi / 3). If it's before the cosine, then the answer is different because you can only invert a positive cosine. If you're not being quizzed on it though, don't worry about that part.

OpenStudy (michele_laino):

By definition, we have: \[{\cos ^{ - 1}}\left( {\cos \left( {\frac{{4\pi }}{5}} \right)} \right) = \frac{{4\pi }}{5} + 2k\pi ,\quad k \in \mathbb{Z}\]

OpenStudy (jdoe0001):

just to state the obvious by now \(\large \bf { cos^{-1}[cos({\color{red}{ \theta}})]={\color{red}{ \theta}} \\ \quad \\ sin^{-1}[sin({\color{red}{ \theta}})]={\color{red}{ \theta}} \\ \quad \\ tan^{-1}[tan({\color{red}{ \theta}})]={\color{red}{ \theta}} }\)